# A dynamical algorithm to compute hyperbolic Julia sets in polynomial time

### v1

@@ Abstract.
Hyperbolic Julia sets of complex polynomials are known to be computable in polynomial time due to pioneering work of Braverman in 2005 [4]. In this paper, we present an alternative method for establishing poly-time computability of hyperbolic Julia sets, which allows us to establish, via a new algorithm, lower computability of the hyperbolicity locus of polynomials of a fixed degree. We first adapt our recently developed algorithms for the computability of polynomial skew products [3], and then apply a refinement that allows us to establish poly-time computation of hyperbilic Julia sets. Finally, we derive lower computability of the hyperbolicity locus via an adapted lattice/refinement search algorithm. In contrast to [4], our approach is dynamical in nature and does not rely on techniques unique to complex analysis.
@@ 1. Introduction
Roughly speaking, our algorithm can be summarized by the following two steps: Step 1: We apply the techniques of [3] to compute a covering of $J$ by finitely many sufficiently small “reference” boxes, so that every reference box has a non-empty intersection with $J$. This succeeds if $p$ is hyperbolic. We note that Step 1 is solely based on dynamical properties including the invariance of the Julia set and its characterization as the expanding chain recurrent component. Moreover, for any given hyperbolic polynomial $p$, Step 1 needs to be performed only once and thus does not contribute to the computational complexity of the precision of the computation of the Julia set of $p$. Step 2: If Step 1 is successful (which it is if $p$ is hyperbolic), we consider a small test box $B$ with center $z$ of side-length $2^{-N}$ where $z$ is contained in (or sufficiently close to) one of the reference boxes. The goal is to decide in polynomial (in fact, linear) complexity of $N$ whether or not the test box intersects the Julia set. To accomplish this we compute underestimates, respectively overestimates, of $p^{k}(B)$ in terms of balls $B(p^{k}(z),c_{u/o}(p^{k})^{\prime}(z))$ for some constants $c_{u}<c_{o}$ which do not depend on $N$. At each iterate we check if the underestimate of $p^{k}(B)$ contains a reference box in which case the algorithm reports $1$, that is, $J$ intersects the test box $B$, or if the overestimate of $p^{k}(B)$ is disjoint from all reference boxes in which case the algorithm reports 0, that is, $B$ does not intersect $J$. We prove that this algorithm terminates if $p$ is hyperbolic, and that the number of required iterates is bounded by a constant times $N$. One of the ingredients to make Step 2 work is a distortion estimate that is used to control the diameter of the iterates of $B$ in terms of the derivative of $p^{k}$ at $z$. While we work here for simplicity with Koebe’s distortion theorem, it should be noted that similar distortion results are known to exist for many smooth uniformly hyperbolic systems.
@@ 2.5. Distortion lemmas
(2.6) $\frac{1}{2}\text{diam}(g(B^{e}(w_{0},\sigma))\geq|g^{\prime}(z_{0})|_{e}\ \sigma\ c(\gamma).$
@@ 3. Results
Moreover, the goal is that the value of $h_{J_{p}}(N,\cdot)$ can be determined (for all $N$ and all input numbers which are dyadic rationals with denominators $2^{N+2}$) via a Turing machine (a computer program for our purposes) whose running time runs is bounded by polynomial function in $N$.
In fact, our algorithm for poly-time computability of $J_{p}$ for a hyperbolic $p$ has a running time which is a constant plus a linear function of $N$. This constant will be an initial, “overhead”, computation coming from applying parts of the algorithms of [3], adapted to a polynomial $p$. The algorithms and results of [3] for polynomial skew products $f(z,w)=(p(z),q(z,w))$ of $\mathbb{C}^{2}$ can be straightforwardly modified to apply in the simpler setting of polynomial maps $p(z)$ of $\mathbb{C}$. In the one-dimensional setting, there are no saddle sets, only the expanding Julia set $J_{p}$ and a finite number of attracting periodic orbits, $A_{p}$.
@@ 3.1. Overhead/first step computation
In this subsection we build the “first step” of our poly-time algorithm; this step is independent of the desired precision ($2^{N}$-approximation of $J_{p}$). In other words, once the polynomial is identified, this part of the algorithm runs once, and halts if $p$ is hyperbolic. Its results can then be used for any $2^{N}$-approximation of the corresponding Julia set. For readability we split this initial algorithm in several parts. First we build a collection of boxes containing $J$ on which the map is expanding.
@@
There exists a Turing machine which based on input of an oracle of a hyperbolic polynomial $p:\mathbb{C}\to\mathbb{C}$ of degree $d\geq 2$, computes dyadic rationals $\delta>0$, $\lambda^{\prime}>0$, an integer $\nu$, and a finite collection of closed square boxes (with disjoint interiors), $\mathcal{B}^{\prime}_{n}=\{B_{j}\}$, such that (i) the boxes’ union contains $J$, i.e., $\mathcal{B}^{\prime}_{n}=\cup_{j}B_{j}\supset J$, and (ii) $p^{\nu}$ is expanding on the boxes’ union; in particular, $|Dp^{\nu}(z)|>1+\lambda^{\prime}$ for all $z\in\mathcal{B}^{\prime}_{n}$.
Further, on $\mathcal{N}_{1}$ defined as the collection of boxes with the same center points as $\mathcal{B}^{\prime}_{n}$ but with radii increased by $2\delta$ (so $\mathcal{N}_{1}:=\bigcup_{B\in\mathcal{B}^{\prime}_{n}}\mathcal{N}(B,2\delta)=\mathcal{N}(\mathcal{B}^{\prime}_{n},2\delta)\supset\cup_{j}B_{j}\supset J$), we still have $p^{\nu}$ expanding: $|Dp^{\nu}(z)|>1+\lambda^{\prime}/2$ for all $z\in\mathcal{N}_{1}$.
@@
We prove by providing the algorithm. First, Algorithm 3.2 of [3] can be applied to $p$ to produce a “box chain model” which includes a set $\mathcal{B}_{n}$ which is a collection of boxes, $\mathcal{B}_{n}=\cup_{j}B_{j}$, all of the same side-length $\varepsilon_{n}$, with pairwise disjoint interiors; in fact, these boxes are a subset of the collection of boxes formed from a $2^{n}\times 2^{n}$ grid on a large box $R,R]^{2}$, where $R$ is a power of $2$ and is calculated to be guaranteed to contain the chain recurrent set $\mathcal{R}$ of $p$ (and hence also contains $J=J_{p}$ and the attracting periodic points $A_{p}$ of $p$). The “model” is this collection of boxes together with a graph $\Gamma_{n}$ whose vertices are the boxes $B_{j}$ and there is guaranteed to be an edge from $B_{k}$ to $B_{j}$ if $p(B_{k})$ intersects $B_{j}$ (but there is guaranteed not to be an edge if $p(B_{k})$ is some given distance away from $B_{j}$); moreover, every box in $\Gamma_{n}$ lies in a cycle. Further, decomposing $\Gamma_{n}$ into strongly connected components yields “box chain components” such that if the boxes are sufficiently small, $J$ and each attracting periodic orbit are contained in different box chain components. (There are no edges in $\Gamma$ connecting different box chain components.)
@@
Algorithm 3.8 (which builds on Algorithm 3.2) of [3], can be adapted to the polynomial map $p$ to compute a sufficiently large $n\in\mathbb{bN}$ and a positive integer $\nu$ such that the box chain model $\mathcal{B}_{n}$ consists entirely of box chain components $\mathcal{B}_{n,i}$ with each determined to be of “expanding” or “contracting” type, and this algorithm halts if $p$ is hyperbolic; specifically, there exist $\lambda_{i}>0$ such that either (a) $|Dp^{\nu}(z)|>1+\lambda_{i}$ for all $z$ in all boxes in $\mathcal{B}_{n,i}$, or (b) $|Dp^{\nu}(z)|<1-\lambda_{i}$ for all $z$ in all boxes in $\mathcal{B}_{n,i}$.
@@
Let $\mathcal{B}^{\prime}_{n}$ denote the collection of all boxes determined to be of expanding type, where we have $|Dp^{\nu}(z)|>1+\lambda^{\prime},$ $\forall z\in\mathcal{B}_{n}^{\prime}$ ($\lambda^{\prime}$ is the minimum of the $\lambda_{i}$’s for all expanding-type components $\mathcal{B}_{n,i}$). Hence we know if $p$ is hyperbolic, $\mathcal{B}^{\prime}_{n}\supset J$.
@@
By applying an argument involving the computable modulus of continuity of the derivative of $p^{\nu}$, as was done similarly in [3], we can compute a bound for the larger neighborhood of $\mathcal{B}^{\prime}_{n}$ on which expansion is only $1+\lambda^{\prime}/2$. In particular, we can compute a lower bound $\delta>0$, and make it a bit smaller so that $\delta$ is a dyadic rational, such that if we replace each box $B$ in $\mathcal{B}^{\prime}_{n}$ by the slightly larger box which is its $2\delta$-neighborhood $\mathcal{N}(B,2\delta)$ in our $L^{\infty}$ metric, we have a collection of boxes with the same center points but with side-lengths larger by $4\delta$ (so the larger box side-length is $4\delta+\varepsilon_{n}$ where $\varepsilon_{n}$ is the box side-length of $\mathcal{B}^{\prime}_{n}$), and we can determine $\delta$ so that all points in these larger boxes satisfy an inequality of expansion but with only a guarantee of weaker expansion $L:=1+\lambda^{\prime}/{2}$; i.e., $|Dp^{\nu}(z)|>1+\lambda^{\prime}/2=L$ for all $z\in\mathcal{N}(\mathcal{B}^{\prime}_{n},2\delta)$. These larger boxes may overlap, but that is no problem. (Using $2\delta$ here is simply convenient for later, $\delta$ will be significant.)
@@
Set $\mathcal{N}_{1}:=\bigcup_{B_{j}\in\mathcal{B}^{\prime}_{n}}\mathcal{N}(B_{j},2\delta)$, so the Hausdorff distance of $\mathcal{B}_{n}^{\prime}$ and $\mathcal{N}_{1}$ is $2\delta$. Since $J\subset\mathcal{B}^{\prime}_{n}$, it follows that $J\subset\mathcal{N}_{1}$ which completes the proof of the lemma. ∎
@@
Given a hyperbolic polynomial $p:\mathbb{C}\to\mathbb{C}$ of degree $d\geq 2$, let $L>1,\delta>0$, $\nu\in\mathbb{bN}$ and $\mathcal{N}_{1}$ denote the finite collection of boxes produced by the Turing machine in Lemma 3.1 so that $|Dp^{\nu}(z)|>L$ for all $z\in\mathcal{N}_{1}\supset\mathcal{N}(J,2\delta).$
There exists a Turing machine which on input of an oracle of a real number $b>0$ computes a finite collection of closed square boxes, $\mathcal{N}_{2}=\cup B_{j}$, of side-length $\beta^{\prime}\in(0,b]$, such that
@@
    the union of the boxes contains the Julia set: $J_{p}\subset\mathcal{N}_{2}$,
@@
    each box contains a point of $J_{p}$: $B_{j}\cap J_{p}\neq\emptyset$ for all $B_{j}\in\mathcal{N}_{2}$,
@@
    the Hausdorff distance from $\mathcal{N}_{2}$ to $\mathcal{N}_{1}$ is at least $\delta$: $\mathcal{N}(\mathcal{N}_{2},\delta)\subset\mathcal{N}_{1}$.
@@
To prove this lemma, we supply the algorithm. As stated the first step is to invoke Lemma 3.1, to create the sets $\mathcal{B}^{\prime}_{n},\mathcal{N}_{1}$, and find $L,\lambda^{\prime},\delta$ and $\nu$. Recall from the proof of the lemma, $\mathcal{B}^{\prime}_{n}$ is a level $n$ box chain model of $p$.
@@
Further, the boxes $\mathcal{B}_{n}$ are contained in $\alpha_{n}$ pseudo-orbits, and an upper bound for $\alpha_{n}$ is computable which converges to 0 as $n\to\infty$. (Theorem 3.6 of [3] includes as a consequence of Algorithm 3.2 of [3] that there is a computable sequence $\alpha_{n}\downarrow 0$ as $n\to\infty$ such that $\mathcal{R}\subset\mathcal{B}_{n}\subset\mathcal{R}(\alpha_{n})$.) Also, the box side-length $\varepsilon_{n}$ (for boxes in $\mathcal{B}^{\prime}_{n}$) satisfies $\varepsilon_{n}<\alpha_{n}$.
We now use a modified part of Algorithm 3.10 of [3]–changed slightly more than just adapting to the one-dimensional case. Here, we describe the outline for this step. We focus on the “(a) Type $J_{2}$-algorithm” part of Algorithm 3.10, which locates a neighborhood of the Julia set (closure of repelling periodic points) to a given precision. Rather than applying this to yield the precision desired based on $N$, what we need is only to subdivide the boxes in $\mathcal{B}^{\prime}_{n}$ just enough to make smaller boxes such that given the expansion $1+\lambda^{\prime}$, the pseudo-orbits in these smaller boxes are $\beta$-shadowed by true orbits for some $\beta>0$. Then for boxes with radius $\beta$ larger, these slightly larger boxes with the same center point each contains a point of $J$, a true orbit. Finally, we need the larger boxes to still be in a neighborhood of $J$ in which $p$ is expanding. This creates an upper bound on $\beta$ (as does the provided $b$). These larger boxes are our set $\mathcal{N}_{2}$. We provide more details in the appendix. ∎
Combining these results produces nested sets with $J\subset\mathcal{N}_{2}\subset\mathcal{N}_{1}$ and the following properties:
@@
  • $|Dg|>1+\lambda^{\prime}/2=L$ on $\mathcal{N}_{2}\subset\mathcal{N}_{1}$,
  • $d_{H}(J,\mathcal{N}_{2})<\beta^{\prime}$, where $\beta^{\prime}$ is the side-lengths of the boxes in $\mathcal{N}_{2}$,
  • $d_{H}(\mathcal{N}_{2},\mathcal{N}_{1})>\delta$, and $d_{H}(\mathcal{N}_{1},J)\geq 2\delta$.

### v2

@@ Abstract.
Hyperbolic Julia sets of complex polynomials are known to be computable in polynomial time due to pioneering work of Braverman in 2005 [4]. In this paper, we present an alternative method for establishing poly-time computability of hyperbolic Julia sets, which allows us to establish, via a new algorithm, lower computability of the hyperbolicity locus of polynomials of a fixed degree. We first adapt our recently developed algorithms for the computability of polynomial skew products [3], and then apply a refinement that allows us to establish poly-time computation of hyperbolic Julia sets. Finally, we derive lower computability of the hyperbolicity locus via an adapted lattice/refinement search algorithm. In contrast to [4], our approach is dynamical in nature and does not rely on techniques unique to complex analysis.
@@ 1. Introduction
Roughly speaking, our algorithm can be summarized by the following two steps. Step 1: We apply the techniques of [3] to compute a covering of $J$ by finitely many sufficiently small “reference” boxes, so that every reference box has a non-empty intersection with $J$. This succeeds if $p$ is hyperbolic. We note that Step 1 is solely based on dynamical properties including the invariance of the Julia set and its characterization as the expanding chain recurrent component. Moreover, for any given hyperbolic polynomial $p$, Step 1 needs to be performed only once and thus does not contribute to the computational complexity of the precision of the computation of the Julia set of $p$. Step 2: If Step 1 is successful (which it is if $p$ is hyperbolic), we consider a small test box $B$ of side-length $2^{-N}$, with center $z$ where $z$ is contained in (or sufficiently close to) one of the reference boxes. The goal is to decide in polynomial (in fact, linear) complexity of $N$ whether or not the test box intersects the Julia set. To accomplish this we compute underestimates, respectively overestimates, of $p^{k}(B)$ in terms of balls $B(p^{k}(z),c_{u/o}(p^{k})^{\prime}(z))$ for some constants $c_{u}<c_{o}$ which are independent of $N$. At each iterate we check if the underestimate of $p^{k}(B)$ contains a reference box, in which case the algorithm reports $1$, that is, $J$ intersects the test box $B$, or if the overestimate of $p^{k}(B)$ is disjoint from all reference boxes, in which case the algorithm reports 0, that is, $B$ does not intersect $J$. We prove that this algorithm terminates if $p$ is hyperbolic, and that the number of required iterates is bounded by a constant times $N$. One of the ingredients to make Step 2 work is a distortion estimate that is used to control the diameter of the iterates of $B$ in terms of the derivative of $p^{k}$ at $z$. While we work here for simplicity with Koebe’s distortion theorem, it should be noted that similar distortion results are known to exist for many smooth uniformly hyperbolic systems.
@@ 2.5. Distortion lemmas
(2.6) $\frac{1}{2}\text{diam}(g(B^{e}(w_{0},\sigma))\geq|g^{\prime}(z_{0})|_{e}\cdot\sigma\cdot c(\gamma).$
@@ 3. Results
Moreover, the goal is that the value of $h_{J_{p}}(N,\cdot)$ can be determined (for all $N$ and all input numbers which are dyadic rationals with denominators $2^{N+2}$) via a Turing machine (a computer program for our purposes) whose running time is bounded by a polynomial function in $N$.
We briefly recall the fundamental difference between our algorithm and Braverman’s approach in [4] (see the discussion after Theorem 1.1 for more details). The algorithm [4] relies on complex-analytic tools—most notably the Poincaré metric and Pick’s theorem—to compute a neighborhood of the Julia set in which points outside the Julia set escape exponentially fast in the Poincaré metric. This escape rate is then algorithmically identified and subsequently translated into escape in the Euclidean metric. Such an approach inherently depends on the holomorphic structure and does not extend to non-holomorphic systems.
In contrast, our method is based entirely on dynamical tools, specifically quantitative shadowing and chain recurrence. We compute a collection of small reference boxes that cover the Julia set, each of which has nonempty intersection with the Julia set. Moreover, we determine a constant $L>1$ such that a suitable (computable) iterate of the polynomial has derivative greater than $L$ in the Euclidean norm throughout each box. Using these expansion bounds together with quantitative escape and distortion estimates, we obtain a criterion to decide whether an arbitrarily small ideal ball contains a point of the Julia set.
We note that our algorithm for poly-time computability of $J_{p}$ for a hyperbolic $p$ has a running time which is a constant plus a linear function of $N$. This constant will be an initial “overhead” computation coming from applying parts of the algorithms of [3], adapted to a polynomial $p$. The algorithms and results of [3] for polynomial skew products $f(z,w)=(p(z),q(z,w))$ of $\mathbb{C}^{2}$ can be straightforwardly modified to apply in the simpler setting of polynomial maps $p(z)$ of $\mathbb{C}$. In the one-dimensional setting, there are no saddle sets, only the expanding Julia set $J_{p}$ and a finite number of attracting periodic orbits, $A_{p}$.
@@ 3.1. Overhead/first step computation
In this subsection, we build the “first step” of our poly-time algorithm. This step is independent of the desired precision ($2^{N}$-approximation of $J_{p}$). In other words, once the polynomial is identified, this part of the algorithm runs once, and halts if $p$ is hyperbolic. Its results can then be used for the $2^{N}$-approximation of the corresponding Julia set, for any $N$. For readability, we split this initial algorithm in several parts. First, we show how to build a collection of boxes containing $J$ on which the map is expanding.
@@
There exists a Turing machine which based on input of an oracle of a hyperbolic polynomial $p:\mathbb{C}\to\mathbb{C}$ of degree $d\geq 2$, computes positive dyadic rationals $\delta$ and $\lambda^{\prime}$, a positive integer $\nu$, and a finite collection of closed square boxes $\mathscr{B}^{\prime}_{n}=\{B_{j}\}$ (with disjoint interiors, which are a subset of the boxes in a $2^{n}\times 2^{n}$ grid on a large box $R,R]^{2}$ containing $\mathcal{R}(p)$), such that:
(i) the boxes’ union contains $J$, i.e., $\mathcal{B}^{\prime}_{n}=\cup_{j}B_{j}\supset J$, and (ii) $p^{\nu}$ is expanding on the boxes’ union; in particular, $|Dp^{\nu}(z)|>1+\lambda^{\prime}$ for all $z\in\mathcal{B}^{\prime}_{n}$.
(iii) Further, on $\mathcal{N}_{1}$ defined as the collection of boxes with the same center points as $\mathcal{B}^{\prime}_{n}$ but with radii increased by $2\delta$ (i.e., $\mathcal{N}_{1}:=\bigcup_{B\in\mathscr{B}^{\prime}_{n}}\mathcal{N}(B,2\delta)=\mathcal{N}(\mathcal{B}^{\prime}_{n},2\delta)$, thus $\mathcal{N}_{1}\supset\mathcal{B}_{n}^{\prime}\supset J$), we have $p^{\nu}$ is still expanding, specifically: $|Dp^{\nu}(z)|>1+\lambda^{\prime}/2$ for all $z\in\mathcal{N}_{1}$.
@@
We prove by providing the algorithm. First, Algorithm 3.2 of [3] can be applied to $p$ to produce a “box chain model” which includes a set $\mathcal{B}_{n}$ which is the union of a collection of boxes, $\mathscr{B}_{n}=\{B_{j}\}$, with the boxes all of the same side-length $\varepsilon_{n}$, with pairwise disjoint interiors; in fact, these boxes are a subset of the collection of boxes formed from a $2^{n}\times 2^{n}$ grid on a large box $R,R]^{2}$, where $R$ is a power of $2$ and is calculated to be guaranteed to contain the chain recurrent set $\mathcal{R}$ of $p$, and hence also contains $J=J_{p}$ and the attracting periodic points $A_{p}$ of $p$. The “model” is this collection of boxes together with a graph $\Gamma_{n}$ whose vertices are the boxes $B_{j}$ and there is guaranteed to be an edge from $B_{k}$ to $B_{j}$ if $p(B_{k})$ intersects $B_{j}$ (and there is guaranteed not to be an edge if $p(B_{k})$ is some given distance away from $B_{j}$); moreover, every box in $\Gamma_{n}$ lies in a cycle. Further, decomposing $\Gamma_{n}$ into strongly connected components yields “box chain components” $\mathcal{B}_{n,i}$ such that if the boxes are sufficiently small, $J$ and each attracting periodic orbit are contained in different box chain components: there are no edges in $\Gamma$ connecting different box chain components.
@@
Algorithm 3.8 (which builds on Algorithm 3.2) of [3], can be adapted to the polynomial map $p$ to compute a sufficiently large $n\in\mathbb{bN}$ and a positive integer $\nu$ such that the box chain model $\mathcal{B}_{n}$ consists entirely of box chain components $\mathcal{B}_{n,i}$ with each determined to be of “expanding” or “contracting” type (for polynomials there is no saddle type), and this algorithm halts if $p$ is hyperbolic; specifically, there exist $\lambda_{i}>0$ such that either (a) $|Dp^{\nu}(z)|>1+\lambda_{i}$ for all $z$ in all boxes in $\mathcal{B}_{n,i}$, or (b) $|Dp^{\nu}(z)|<1-\lambda_{i}$ for all $z$ in all boxes in $\mathcal{B}_{n,i}$.
@@
We let $\mathscr{B}^{\prime}_{n}$ denote the collection of all boxes determined to be of expanding type, and we let $\mathcal{B}^{\prime}_{n}$ be their union. We have $|Dp^{\nu}(z)|>1+\lambda^{\prime},$ $\forall z\in\mathcal{B}_{n}^{\prime}$ (where $\lambda^{\prime}$ is the minimum of the $\lambda_{i}$’s for all expanding-type components $\mathcal{B}_{n,i}$). Hence we know if $p$ is hyperbolic, $\mathcal{B}^{\prime}_{n}\supset J$.
@@
By applying an argument involving the computable modulus of continuity of the derivative of $p^{\nu}$, as was done similarly in [3], we can compute a bound for a larger neighborhood of $\mathcal{B}^{\prime}_{n}$ on which expansion is only $1+\lambda^{\prime}/2$. In particular, we can compute a lower bound $\delta>0$, and make it a bit smaller so that $\delta$ is a dyadic rational, such that if we replace each box $B$ in $\mathscr{B}^{\prime}_{n}$ by the slightly larger box which is its $2\delta$-neighborhood $\mathcal{N}(B,2\delta)$ in our $L^{\infty}$ metric (using $2\delta$ here is convenient as later, $\delta$ will be significant), we have a collection of boxes with the same center points but with side-lengths larger by $4\delta$ (so the larger box side-length is $4\delta+\varepsilon_{n}$ where $\varepsilon_{n}$ is the box side-length of $\mathcal{B}^{\prime}_{n}$), such that all points in these larger boxes satisfy an inequality of expansion but only with a guarantee of weaker expansion $L:=1+\lambda^{\prime}/{2}$. That is, $|Dp^{\nu}(z)|>1+\lambda^{\prime}/2=L$ for all $z\in\mathcal{N}(\mathcal{B}^{\prime}_{n},2\delta)$. These larger boxes may overlap, but that is no problem.
@@
Set $\mathcal{N}_{1}:=\bigcup_{B_{j}\in\mathscr{B}^{\prime}_{n}}\mathcal{N}(B_{j},2\delta)$, so the Hausdorff distance of $\mathcal{B}_{n}^{\prime}$ and $\mathcal{N}_{1}$ is exactly $2\delta$ (i.e., $d(H(\mathcal{B}^{\prime}_{n},\mathcal{N}_{1})=2\delta$). Since $J\subset\mathcal{B}^{\prime}_{n}$, it follows that $J\subset\mathcal{N}_{1}$, which completes the proof of the lemma. Note we also have $d_{H}(J,\mathcal{N}_{1})\geq 2\delta$ and $\mathcal{N}_{1}\supset\mathcal{N}(J,2\delta).$
@@
Given a hyperbolic polynomial $p:\mathbb{C}\to\mathbb{C}$ of degree $d\geq 2$, suppose $L>1,\delta>0$, $\nu\in\mathbb{bN}^{+}$ and $\mathcal{N}_{1}$ are produced by the Turing machine in Lemma 3.1. So, we have $|Dp^{\nu}(z)|>L$ for all $z\in\mathcal{N}_{1}\supset\mathcal{N}(J,2\delta).$
Then there exists a Turing machine which on input of an oracle of a real number $b>0$ computes a set which is a finite union of closed square boxes, $\mathcal{N}_{2}=\cup B_{j}$, with each box of side-length $\beta^{\prime}\in(0,b]$, such that
@@
    the union of these boxes contains the Julia set: $J_{p}\subset\mathcal{N}_{2}$,
@@
    each box $B_{j}$ which is an element of the union of boxes defining $\mathcal{N}_{2}$ contains a point of $J_{p}$: $B_{j}\cap J_{p}\neq\emptyset$ for all $B_{j}$ in $\cup B_{j}=\mathcal{N}_{2}$,
@@
    the Hausdorff distance from $\mathcal{N}_{2}$ to $\mathcal{N}_{1}$ is at least $\delta$, and a $\delta$-neighborhood of $\mathcal{N}_{2}$ is contained in $\mathcal{N}_{1}$: $d_{H}(\mathcal{N}_{2},\mathcal{N}_{1})>\delta$ and $\mathcal{N}(\mathcal{N}_{2},\delta)\subset\mathcal{N}_{1}$.
We call these boxes “reference boxes”.
@@
To prove this lemma, we supply the algorithm. The first step is to invoke Lemma 3.1, creating the sets $\mathcal{B}^{\prime}_{n},\mathcal{N}_{1}$, and finding the constants $L,\lambda^{\prime},\delta$ and $\nu$. Recall from the proof of the lemma, $\mathcal{B}^{\prime}_{n}$ is a level $n$ box chain model of $p$, and we have $|Dp^{\nu}(z)|>L=1+\lambda^{\prime}/2$ for all $z\in\mathcal{N}_{1}\supset\mathcal{N}(J,2\delta).$
@@
Further, Theorem 3.6 of [3] includes (as a consequence of Algorithm 3.2 of [3]) that there is a computable sequence $\alpha_{n}\downarrow 0$ as $n\to\infty$ such that $\mathcal{R}\subset\mathcal{B}_{n}\subset\mathcal{R}(\alpha_{n})$. Hence, the boxes defining $\mathcal{B}_{n}$ are contained in $\alpha_{n}$-pseudo-orbits, and an upper bound for $\alpha_{n}$ is computable and converges to 0 as $n\to\infty$. Also, the box side-length $\varepsilon_{n}$ (for boxes defining $\mathcal{B}^{\prime}_{n}$) satisfies $\varepsilon_{n}<\alpha_{n}$.
We now use a modified part of Algorithm 3.10 of [3]–changed slightly more than just adapting to the one-dimensional case. Here, we describe the outline for this step. We focus on the “(a) Type $J_{2}$-algorithm” part of Algorithm 3.10, which locates a neighborhood of the Julia set (closure of repelling periodic points) to a given precision. Rather than applying this to yield the precision desired based on $N$, what we need is only to subdivide the boxes in $\mathcal{B}^{\prime}_{n}$ just enough to make smaller boxes such that given the expansion $1+\lambda^{\prime}$, the pseudo-orbits in these smaller boxes are $\beta$-shadowed by true orbits for some $\beta>0$. Then for boxes with radius $\beta$ larger, these slightly larger boxes with the same center point each contains a point of $J$, a true orbit; so the sidelength of these boxes which each contain a point of $J$ is our $\beta^{\prime}$. Finally, we need the $\beta^{\prime}$-boxes to still be in the neighborhood of $J$ in which $p$ is expanding (that is, $\mathcal{N}_{1}$). This creates another upper bound on $\beta^{\prime}$ (as does the provided $b$). The union of these $\beta^{\prime}$-sidelength boxes defines our set $\mathcal{N}_{2}$. We provide more details in the appendix, establishing items (i)–(iv). ∎
Combining these results produces nested sets with $J\subset\mathcal{N}_{2}\subset\mathcal{N}_{1}$ and the following properties for $g=p^{\nu}$:
@@
  • $|Dg|>1+\lambda^{\prime}/2=L$ on both $\mathcal{N}_{2}$ and $\mathcal{N}_{1}$,
  • $d_{H}(\mathcal{N}_{1},J)\geq 2\delta$ (since $d_{H}(\mathcal{B}^{\prime}_{n},\mathcal{N}_{1})=2\delta$ and $J\subset\mathcal{B}^{\prime}_{n}$) and $\mathcal{N}_{1}\supset\mathcal{N}(J,2\delta),$
  • $d_{H}(J,\mathcal{N}_{2})<\beta^{\prime}$, where $\beta^{\prime}$ is the side-lengths of the boxes in $\mathcal{N}_{2}$, and
  • $d_{H}(\mathcal{N}_{2},\mathcal{N}_{1})>\delta$ and $\mathcal{N}(\mathcal{N}_{2},\delta)\subset\mathcal{N}_{1}$.